find the discriminant of each quadratic equation.
A. x^ + 4x +9 =0
B.2x^ + 3x =0

Answer by stanbon(75887)

find the discriminant of each quadratic equation.
A. x^2 + 4x +9 =0
b^2-4ac = 16-4*1*9 = -20
---------------------------------
B.2x^2 + 3x =0
b^2-4ac = 9-4*2*0 = 9
========================
Cheers,
stan N